OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form? x – 4y = –18 x + 4y = 18 4x – y = –18 4x + y = 18

OpenStudy (anonymous):

let the line be Y=MX+C as its passing through(-3,6) so 6=-3M+C given also its parallel to line y=4x-2 and we know that 2 lines parallel have product of their slopes as -1 i.e. M*(4)=-1 or M=-1/4 so 6=-3*(-1/4)+C C=6-3/4 C=21/4 hence equation of desired line is Y=(-1/4)X+21/4 X+4Y=21 i guess it option B

OpenStudy (anonymous):

the line you are looking for is parallel so the slope will be the same as the original line you were given, which is slope of 4 Now we will use point slope form because we have points and a slope. y - y1 = m(x - x1) slope(m) = 4 (-3,6) x1 = -3, y1 = 6 y - 6 = 4(x - (-3) y - 6 = 4(x + 3) y - 6 = 4x + 12 y = 4x + 12 + 6 y = 4x + 18 standard form would be : -4x + y = 18 (multiply by -1) 4x - y = -18 (divide by 4) x - 1/4y= -18/4 I am getting something totally different ???

OpenStudy (anonymous):

no wait....standard form is 4x - y = -18

OpenStudy (anonymous):

so which one is the right answer ?

OpenStudy (anonymous):

you go with who you want....but I am saying it is 4x - y = -18

OpenStudy (anonymous):

or wait for somebody else to come and see what they get

OpenStudy (anonymous):

I'll try that one i appriciate it ! this is like a time test n i need to pass my algebra 2 to graduate im strugglinf thank youu

OpenStudy (anonymous):

your welcome :)